7. A worker's total earnings for one day is $100. He received a $20 fixed payment and





consumes 14 hours of leisure. What is the hourly wage rate? Show solutions

1
Expert's answer
2022-02-15T11:21:26-0500

SOLUTION


Total earnings = fixed earnings + hourly earnings


Assuming hourly wage rate to be h


100 = 20 + h (24 - 14)

100 = 20 + 10h

100 - 20 = 10h

80 = 10h

8 = h


ANSWER

The hourly wage rate is $8